# AdaptVision: Efficient Vision-Language Models via Adaptive Visual Acquisition

### v1

@@ 1 Introduction
Recently, Vision-Language Models (VLMs) [li2023blip, bai2023qwen, chen2024sharegpt4v] have achieved significant breakthroughs in general visual question answering (VQA) and diverse practical applications by projecting and adapting visual tokens into large language model (LLM) space [touvron2023llama, achiam2023gpt, zhu2023minigpt, bai2023qwen]. However, the promising performance of VLMs largely relies on the large amount of vision tokens, inevitably introducing a huge memory and computational overhead when compared to LLMs, particularly for high-resolution images. For instance, a $2048\times 1024$ image yields 2,678 vision tokens in Qwen2.5-VL [bai2025qwen2]. Therefore, it is crucial to avoid the excessive consumption of visual tokens.
Numerous studies have explored visual token compression to enhance VLM efficiency [yang2025visionzip, chen2024image, wen2024efficient, shi2023upop, he2024zipvl, jian2023expedited, zhang2024sparsevlm, yang2025visionthink]. Existing works can be categorized into two main research directions. The first prunes or merges a fixed number of visual tokens based on predetermined thresholds, according to the importance and similarity of vision tokens [yang2025visionzip, chen2024image, zhang2024sparsevlm]. The second dynamically processes distinct samples, where the system adaptively switches between using 100% vision tokens for OCR-related tasks and 25% vision tokens for simpler tasks by selectively employing quarter-resolution images [yang2025visionthink]. However, existing efficient VLM paradigms and methods are largely passive, as they can only reduce the number of vision tokens by predefined ratios. This leads to a natural question: Can VLMs adaptively determine the minimum number of vision tokens for each sample according to different scenarios?
Cognitive neuroscience reveals that our visual system operates through an active, sequential, and adaptive process known as active vision [findlay2003active, itti2005neurobiology]. It first captures coarse, low-spatial-frequency information to grasp the gist of a scene, then directs attention to salient regions for detailed analysis [navon1977forest]. This coarse-to-fine processing mechanism enables humans to efficiently parse complex visual inputs with minimal cognitive load. Fig. 1 provides an illustrative example.
The cognitive strategy of active vision is operationalized in recent VLMs through the “thinking-with-images” paradigm, such as invoking tools to zoom and crop specific regions [zheng2025deepeyes, lai2025mini] to advance fine-grained visual understanding. We argue that this active reasoning capability can be effectively applied to the critical task of visual token reduction, allowing the model to decide how few visual tokens are sufficient.
@@
However, training this dual-objective policy with standard RL algorithms like Group Relative Policy Optimization (GRPO) [shao2024deepseekmath] presents two key challenges: (1) Ambiguous credit assignment: Vanilla GRPO assigns a single sequence-level reward to all generated tokens, failing to distinguish the contribution of the decision to request additional visual tokens from that of generating the final answer; (2) Imbalanced optimization: Since vanilla GRPO normalizes all tokens uniformly in a sequence, it introduces an imbalance: compared to 1-turn direct-answer sequences, 2-turn tool-invoking sequences receive imbalanced gradient signals, causing the latter to be under-optimized.
@@ 2 Related work
Recent advances in reasoning LLMs such as OpenAI’s o1 [jaech2024openai] and DeepSeek R1 [guo2025deepseek] have accelerated the use of RL to enhance reasoning capabilities. This trend has extended to VLMs [tan2025reason, shen2025vlm, liu2025visual, peng2025lmm], where most work focuses on high-level semantic reasoning like tool use or chain-of-thought explanation. A related direction explores active perception, equipping VLMs with fine-grained control mechanisms [wu2024v, huang2025high, su2025pixel]. Recent systems such as DeepEyes [zheng2025deepeyes] and Mini-o3 [lai2025mini] support operations like zoom and crop, improving performance on detailed visual tasks. While these methods showcase the power of active visual reasoning for enhancing answer accuracy, how to apply this “thinking with images” paradigm to the goal of computational efficiency, specifically for visual token compression, remains less explored. Our method enables the VLM to autonomously determine the minimum number of visual tokens required for a given task, thereby achieving efficient inference while maintaining performance.
@@
Reducing VLM computational cost by vision token compression has become a popular research topic. Existing methods rely on predefined rules or metrics to compress tokens. For instance, FastV [chen2024image] prunes a fixed 50% of tokens based on attention scores after the second layer. PyramidDrop [xing2024pyramiddrop] proposes progressive token compression to reduce information loss. Other works leverage cross-modal relevance for token selection, such as SparseVLM [zhang2024sparsevlm] and VisionZip [yang2025visionzip], which retain semantically relevant visual tokens. A key limitation of these methods is their dependence on a fixed compression ratio, which lacks adaptability across tasks. VisionThink [yang2025visionthink] uses RL to decide whether to use a low-resolution or the original image, offering limited adaptability but still restricting the model to coarse-grained decisions. In contrast, our approach enables the VLM to learn coarse-to-fine ability and adaptively determine the minimum number of visual tokens for each task.
@@ 3.1 Reinforcement Learning for LLMs
Recent studies [guo2025deepseek, jaech2024openai] have demonstrated RL effectively enhances the reasoning capabilities of large language models (LLMs). Recently, Group Relative Policy Optimization (GRPO) [shao2024deepseekmath] has been widely used in LLM reasoning. Given a question $x$, GRPO generates $G$ distinct responses $\{o_{i}\}_{i=1}^{G}$ with sequence length $N_{i}$ from the current policy $\pi_{\theta_{old}}$ and obtains a group of rewards $\{R_{i}\}_{i=1}^{G}$. GRPO optimizes the policy model $\pi_{\theta}$ by maximizing the following objective:
@@ 3.2 Vision Language Models
The VLM architectures generally consist of three components: a visual encoder, a modality projector, and a LLM. A commonly used approach for the visual encoder is to employ a pre-trained image encoder like CLIP-VIT [radford2021learning] that converts input images into visual tokens. The modality projector adjusts the size of these visual tokens to match the embedding size of LLM and to achieve semantic alignment, enabling the LLM to process visual data effectively. The LLM then integrates the aligned visual and textual information to generate responses.
Existing works have revealed that the computational complexity of VLM is strongly influenced by the sequence length [yang2025visionzip], where the sequence length is defined as $n=n_{sys}+n_{img}+n_{question}$. In typical VLM tasks, the number of vision tokens $n_{img}$ is often much larger than the other two, sometimes by a factor of 20. Therefore, reducing the number of vision tokens is the key for improving the efficiency of VLMs.
@@ 4.2 Reward Design
The outcome reward is the sum of three components. (1) Accuracy reward $\mathcal{R}_{acc}$: Since VQA answers are typically open-ended, we use an LLM as judge to assign a binary reward (1 for correct, 0 for incorrect) for answer correctness. (2) Format reward $\mathcal{R}_{form}$: To maintain instruction-following capability, we enforce formatting requirements: reasoning in <think> tags, answers in <answer> tags, and tool calls in <tool_call> tags with valid JSON. The format reward is 0.5 for full compliance with all formatting requirements; otherwise, the reward is 0. (3) Balance reward $\mathcal{R}_{bal}$: To prevent over-reliance on tool calls, we introduce a balance reward. We apply a 0.1 penalty to correct answers that invoke tool calls. Additionally, to discourage “lucky guesses” [yang2025visionthink], we impose a 0.1 penalty on direct answers when the probability of correct response from low-resolution images is low, thereby encouraging appropriate tool usage. The design of this balance reward is as follows:
@@ 4.3 Efficient Learning via Decoupled Turn Policy Optimization
Based on our reward design, we initially employ GRPO [shao2024deepseekmath] for training. We aim to train a VLM that (1) achieves high answering accuracy and (2) minimizes the number of visual tokens used. However, training such a dual-objective policy with GRPO presents two key challenges.
@@ 5.1 Evaluation Setup
We conduct experiments on several general VQA benchmarks, including ChartQA [masry2022chartqa], OCRBench [liu2024ocrbench], DocVQA [mathew2021docvqa], MME [fu2024mmecomprehensiveevaluationbenchmark], MMVet [yu2023mm], RealWorldQA [xai2024grok1.5v_online], POPE [li2023evaluatingobjecthallucinationlarge], MathVista [lu2023mathvista], MathVerse [zhang2024mathverse]. AdaptVision is based on Qwen2.5-VL-7B-Instruct [bai2025qwen2]. We employ veRL [sheng2025hybridflow] framework for RL training. During training, we set the batch size as 512 and the mini-batch size as 32. We drop the KL term during policy optimization. The initial learning rate of the policy model is 1e61e-6. For each prompt, we sample 16 candidate responses using a temperature of 1.0. During inference, we use the vLLM framework and set the temperature to 0. We use training data from yang2025visionthink [^1] https://huggingface.co/datasets/Senqiao/VisionThink-Smart-Train, which contains VQA samples that can be answered directly using low-resolution images, as well as samples that require high-resolution images for accurate answering.
@@ 5.2 Main Results
We compare AdaptVision with existing vision token compression methods, including FastV [chen2024image], SparseVLM [zhang2024sparsevlm], VisionZip [yang2025visionzip], and VisionThink [yang2025visionthink]. FastV, SparseVLM, and VisionZip are static methods that operate with a pre-defined token retention ratio, while VisionThink and AdaptVision are dynamic methods that vary visual token usage for each sample. For fair comparison, static methods are set to 50% token retention. For VisionThink, we initially used the officially released model [^2] https://huggingface.co/Senqiao/VisionThink-Efficient but found it consumed substantially more visual tokens than our method, making the comparison unfair. We thus report two versions: “VisionThink” for the released model and “VisionThink” for our reproduction using the public code [^3] https://github.com/dvlab-research/VisionThink. We also include the vanilla model (100% tokens, high-resolution) and the down-sample model (25% tokens, 1/4 resolution) based on Qwen2.5-VL-7B-Instruct. Results are shown in Table 1. Compared to previous vision token compression methods (including FastV, SparseVLM, VisionZip and VisionThink), AdaptVision achieves superior average performance across all benchmarks with significantly fewer visual tokens. Compared to the down-sample model, AdaptVision improves accuracy by 5.8% (92.1% $\rightarrow$ 97.9%) with only 7% (25% $\rightarrow$ 33%) more visual tokens, highlighting its effective coarse-to-fine visual reasoning.
@@ A.1 Prompt Details
AdaptVision utilizes three types of prompts. First, to equip the VLM with basic tool-using capability, we follow the Qwen2.5-VL cookbook [bai2025qwen2] to design prompts for the bounding box tool (Table 3). Second, since VQA tasks are typically diverse and open-ended, we adopt an LLM-as-judge approach to evaluate answer correctness. As shown in Table 4, following yang2025visionthink, we design a judging prompt for GPT-4o to produce binary evaluations (1 for correct, 0 for incorrect). Third, to encourage efficient tool exploration, we prompt GPT-4o to evaluate the relevance of cropped regions, producing a binary reward for region correctness (Table 5).
@@ A.2 Training and Evaluation Details
AdaptVision is based on Qwen2.5-VL-7B-Instruct [bai2025qwen2]. We employ veRL [sheng2025hybridflow] framework for RL training. During training, we set the batch size as 512 with mixed-precision (FP16) training. The mini-batch size is 32. We drop the KL term during policy optimization. For each prompt, we sample 16 candidate responses (i.e., $G=16$) using a temperature of 1.0. The upper and lower clip ratios are 0.24 and 0.20, respectively. We set the maximum prompt length and the maximum response length as 8192. All experiments were conducted on 4 nodes, each with 8 H20 GPUs. The model was trained for 80 steps, using the AdamW optimizer with a learning rate of 1e61e-6, $\beta=(0.9,0.999)$, and a weight decay of 0.01. During inference, we use the vLLM framework and set the temperature to 0.

### v2

# arXiv:2512.03794v2[cs.CV] 28 Feb 2026

@@ 1 Introduction
Recently, Vision-Language Models (VLMs) [15, 2, 5] have achieved significant breakthroughs in general visual question answering (VQA) and diverse practical applications by projecting and adapting visual tokens into large language model (LLM) space [31, 1, 42, 2]. However, the promising performance of VLMs largely relies on the large amount of vision tokens, inevitably introducing a huge memory and computational overhead when compared to LLMs, particularly for high-resolution images. For instance, a $2048\times 1024$ image yields 2,678 vision tokens in Qwen2.5-VL [3]. Therefore, it is crucial to avoid the excessive consumption of visual tokens.
Numerous studies have explored visual token compression to enhance VLM efficiency [36, 4, 32, 28, 9, 13, 40, 37]. Existing works can be categorized into two main research directions. The first prunes or merges a fixed number of visual tokens based on predetermined thresholds, according to the importance and similarity of vision tokens [36, 4, 40]. The second dynamically processes distinct samples, where the system adaptively switches between using 100% vision tokens for OCR-related tasks and 25% vision tokens for simpler tasks by selectively employing quarter-resolution images [37]. However, existing efficient VLM paradigms and methods are largely passive, as they can only reduce the number of vision tokens by predefined ratios. This leads to a natural question: Can VLMs adaptively determine the minimum number of vision tokens for each sample according to different scenarios?
Cognitive neuroscience reveals that our visual system operates through an active, sequential, and adaptive process known as active vision [6, 11]. It first captures coarse, low-spatial-frequency information to grasp the gist of a scene, then directs attention to salient regions for detailed analysis [22]. This coarse-to-fine processing mechanism enables humans to efficiently parse complex visual inputs with minimal cognitive load. Fig. 1 provides an illustrative example.
The cognitive strategy of active vision is operationalized in recent VLMs through the “thinking-with-images” paradigm, such as invoking tools to zoom and crop specific regions [41, 14] to advance fine-grained visual understanding. We argue that this active reasoning capability can be effectively applied to the critical task of visual token reduction, allowing the model to decide how few visual tokens are sufficient.
@@
However, training this dual-objective policy with standard RL algorithms like Group Relative Policy Optimization (GRPO) [25] presents two key challenges: (1) Ambiguous credit assignment: Vanilla GRPO assigns a single sequence-level reward to all generated tokens, failing to distinguish the contribution of the decision to request additional visual tokens from that of generating the final answer; (2) Imbalanced optimization: Since vanilla GRPO normalizes all tokens uniformly in a sequence, it introduces an imbalance: compared to 1-turn direct-answer sequences, 2-turn tool-invoking sequences receive imbalanced gradient signals, causing the latter to be under-optimized.
@@ 2 Related work
Recent advances in reasoning LLMs such as OpenAI’s o1 [12] and DeepSeek R1 [8] have accelerated the use of RL to enhance reasoning capabilities. This trend has extended to VLMs [30, 26, 18, 23], where most work focuses on high-level semantic reasoning like tool use or chain-of-thought explanation. A related direction explores active perception, equipping VLMs with fine-grained control mechanisms [33, 10, 29]. Recent systems such as DeepEyes [41] and Mini-o3 [14] support operations like zoom and crop, improving performance on detailed visual tasks. While these methods showcase the power of active visual reasoning for enhancing answer accuracy, how to apply this “thinking with images” paradigm to the goal of computational efficiency, specifically for visual token compression, remains less explored. Our method enables the VLM to autonomously determine the minimum number of visual tokens required for a given task, thereby achieving efficient inference while maintaining performance.
@@
Reducing VLM computational cost by vision token compression has become a popular research topic. Existing methods rely on predefined rules or metrics to compress tokens. For instance, FastV [4] prunes a fixed 50% of tokens based on attention scores after the second layer. PyramidDrop [35] proposes progressive token compression to reduce information loss. Other works leverage cross-modal relevance for token selection, such as SparseVLM [40] and VisionZip [36], which retain semantically relevant visual tokens. A key limitation of these methods is their dependence on a fixed compression ratio, which lacks adaptability across tasks. VisionThink [37] uses RL to decide whether to use a low-resolution or the original image, offering limited adaptability but still restricting the model to coarse-grained decisions. In contrast, our approach enables the VLM to learn coarse-to-fine ability and adaptively determine the minimum number of visual tokens for each task.
@@ 3.1 Reinforcement Learning for LLMs
Recent studies [8, 12] have demonstrated RL effectively enhances the reasoning capabilities of large language models (LLMs). Recently, Group Relative Policy Optimization (GRPO) [25] has been widely used in LLM reasoning. Given a question $x$, GRPO generates $G$ distinct responses $\{o_{i}\}_{i=1}^{G}$ with sequence length $N_{i}$ from the current policy $\pi_{\theta_{old}}$ and obtains a group of rewards $\{R_{i}\}_{i=1}^{G}$. GRPO optimizes the policy model $\pi_{\theta}$ by maximizing the following objective:
@@ 3.2 Vision Language Models
The VLM architectures generally consist of three components: a visual encoder, a modality projector, and a LLM. A commonly used approach for the visual encoder is to employ a pre-trained image encoder like CLIP-VIT [24] that converts input images into visual tokens. The modality projector adjusts the size of these visual tokens to match the embedding size of LLM and to achieve semantic alignment, enabling the LLM to process visual data effectively. The LLM then integrates the aligned visual and textual information to generate responses.
Existing works have revealed that the computational complexity of VLM is strongly influenced by the sequence length [36], where the sequence length is defined as $n=n_{sys}+n_{img}+n_{question}$. In typical VLM tasks, the number of vision tokens $n_{img}$ is often much larger than the other two, sometimes by a factor of 20. Therefore, reducing the number of vision tokens is the key for improving the efficiency of VLMs.
@@ 4.2 Reward Design
The outcome reward is the sum of three components. (1) Accuracy reward $\mathcal{R}_{acc}$: Since VQA answers are typically open-ended, we use an LLM as judge to assign a binary reward (1 for correct, 0 for incorrect) for answer correctness. (2) Format reward $\mathcal{R}_{form}$: To maintain instruction-following capability, we enforce formatting requirements: reasoning in <think> tags, answers in <answer> tags, and tool calls in <tool_call> tags with valid JSON. The format reward is 0.5 for full compliance with all formatting requirements; otherwise, the reward is 0. (3) Balance reward $\mathcal{R}_{bal}$: To prevent over-reliance on tool calls, we introduce a balance reward. We apply a 0.1 penalty to correct answers that invoke tool calls. Additionally, to discourage “lucky guesses” [37], we impose a 0.1 penalty on direct answers when the probability of correct response from low-resolution images is low, thereby encouraging appropriate tool usage. The design of this balance reward is as follows:
@@ 4.3 Efficient Learning via Decoupled Turn Policy Optimization
Based on our reward design, we initially employ GRPO [25] for training. We aim to train a VLM that (1) achieves high answering accuracy and (2) minimizes the number of visual tokens used. However, training such a dual-objective policy with GRPO presents two key challenges.
@@ 5.1 Evaluation Setup
We conduct experiments on several general VQA benchmarks, including ChartQA [20], OCRBench [17], DocVQA [21], MME [7], MMVet [38], RealWorldQA [34], POPE [16], MathVista [19], MathVerse [39]. AdaptVision is based on Qwen2.5-VL-7B-Instruct [3]. We employ veRL [27] framework for RL training. During training, we set the batch size as 512 and the mini-batch size as 32. We drop the KL term during policy optimization. The initial learning rate of the policy model is 1e61e-6. For each prompt, we sample 16 candidate responses using a temperature of 1.0. During inference, we use the vLLM framework and set the temperature to 0. We use training data from Yang et al. [37] [^1] https://huggingface.co/datasets/Senqiao/VisionThink-Smart-Train, which contains VQA samples that can be answered directly using low-resolution images, as well as samples that require high-resolution images for accurate answering.
@@ 5.2 Main Results
We compare AdaptVision with existing vision token compression methods, including FastV [4], SparseVLM [40], VisionZip [36], and VisionThink [37]. FastV, SparseVLM, and VisionZip are static methods that operate with a pre-defined token retention ratio, while VisionThink and AdaptVision are dynamic methods that vary visual token usage for each sample. For fair comparison, static methods are set to 50% token retention. For VisionThink, we initially used the officially released model [^2] https://huggingface.co/Senqiao/VisionThink-Efficient but found it consumed substantially more visual tokens than our method, making the comparison unfair. We thus report two versions: “VisionThink” for the released model and “VisionThink” for our reproduction using the public code [^3] https://github.com/dvlab-research/VisionThink. We also include the vanilla model (100% tokens, high-resolution) and the down-sample model (25% tokens, 1/4 resolution) based on Qwen2.5-VL-7B-Instruct. Results are shown in Table 1. Compared to previous vision token compression methods (including FastV, SparseVLM, VisionZip and VisionThink), AdaptVision achieves superior average performance across all benchmarks with significantly fewer visual tokens. Compared to the down-sample model, AdaptVision improves accuracy by 5.8% (92.1% $\rightarrow$ 97.9%) with only 7% (25% $\rightarrow$ 33%) more visual tokens, highlighting its effective coarse-to-fine visual reasoning.
@@ A.1 Prompt Details
AdaptVision utilizes three types of prompts. First, to equip the VLM with basic tool-using capability, we follow the Qwen2.5-VL cookbook [3] to design prompts for the bounding box tool (Table 3). Second, since VQA tasks are typically diverse and open-ended, we adopt an LLM-as-judge approach to evaluate answer correctness. As shown in Table 4, following Yang et al. [37], we design a judging prompt for GPT-4o to produce binary evaluations (1 for correct, 0 for incorrect). Third, to encourage efficient tool exploration, we prompt GPT-4o to evaluate the relevance of cropped regions, producing a binary reward for region correctness (Table 5).
@@ A.2 Training and Evaluation Details
AdaptVision is based on Qwen2.5-VL-7B-Instruct [3]. We employ veRL [27] framework for RL training. During training, we set the batch size as 512 with mixed-precision (FP16) training. The mini-batch size is 32. We drop the KL term during policy optimization. For each prompt, we sample 16 candidate responses (i.e., $G=16$) using a temperature of 1.0. The upper and lower clip ratios are 0.24 and 0.20, respectively. We set the maximum prompt length and the maximum response length as 8192. All experiments were conducted on 4 nodes, each with 8 H20 GPUs. The model was trained for 80 steps, using the AdamW optimizer with a learning rate of 1e61e-6, $\beta=(0.9,0.999)$, and a weight decay of 0.01. During inference, we use the vLLM framework and set the temperature to 0.